OpenStudy (anonymous):

ln(2x)=3^-x+2: solve for x

OpenStudy (anonymous):

Can I help?

OpenStudy (anonymous):

yes please

OpenStudy (anonymous):

help please

OpenStudy (anonymous):

i am here no screeming

OpenStudy (anonymous):

can you help?

OpenStudy (anonymous):

yes afcours

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

whait a second my cancalater glitching bad writing right becouse i am russian ahhaa

OpenStudy (anonymous):

\[\ln (2x)=3^{-x+2}\]

OpenStudy (anonymous):

that makes more sence

OpenStudy (anonymous):

i dont understand it correctly i can do it only on 70%

OpenStudy (anonymous):

i will call for help

OpenStudy (anonymous):

@DanJS @timo86m @jabez177 @tHe_FiZiCx99 @imqwerty @inkyvoyd @Koikkara @Jemurray3 @rishavraj @sweetburger @Yttrium @narissa @MTALHAHASSAN2 @AloneS @Kitten_is_back @Empty

OpenStudy (anonymous):

what this for ?

OpenStudy (anonymous):

when you need help you call people like that

OpenStudy (anonymous):

are they good ?

OpenStudy (anonymous):

it says in there mail box some one called there name to help

OpenStudy (anonymous):

yes they are experts

OpenStudy (anonymous):

email them for help?

OpenStudy (anonymous):

i did

OpenStudy (anonymous):

just scroll around there name and you can see there points DanJS is my favirate one

OpenStudy (anonymous):

ok , I will try

jabez177 (jabez177):

http://prntscr.com/9t90jc

OpenStudy (anonymous):

here is one of them that i called

jabez177 (jabez177):

Lol... I'm doing Physics right now so I'm gone. :P :)

OpenStudy (anonymous):

hi @jabez177

OpenStudy (anonymous):

can you hep?

OpenStudy (anonymous):

how me

OpenStudy (anonymous):

any one can help

OpenStudy (anonymous):

but he gave you the answere

OpenStudy (anonymous):

who?, no one answer yes

OpenStudy (anonymous):

the answere is 0 once again

OpenStudy (anonymous):

no, the key is 1.78

OpenStudy (inkyvoyd):

hmm

OpenStudy (inkyvoyd):

ln(2x)=3^(-x+2) right?

OpenStudy (anonymous):

right

OpenStudy (anonymous):

where did you come from

OpenStudy (dumbcow):

this cant be solved algebraically, because you need equation to be all logs or all exponential not mixed

OpenStudy (anonymous):

o ya i forgot i called you

OpenStudy (anonymous):

I know that why hard to solve

OpenStudy (anonymous):

haaaaaaaaaaaaaaaaaaaaaaaaaaaaaaaaaaaaaaaaaaaaaaaaaaaaaaaaaaaaaaaaaaaaaaaaaaaaaaaaaaaaaaaaaaaaaaaaaaaaaaaaaaahahahahah

OpenStudy (anonymous):

inkyvoyd that is a strange que.......... Nancy

OpenStudy (dumbcow):

the answer is from using numerical approximation, im guessing newtons method which involves calculus

OpenStudy (anonymous):

can you solve by step?

OpenStudy (inkyvoyd):

dumbcow is right here... it's too late at night for me to be doing mathematics.

OpenStudy (anonymous):

ok, we will do tomorrow . thanks

OpenStudy (anonymous):

i just woke up it is morning in russia

OpenStudy (anonymous):

dumbcow you write slow

OpenStudy (anonymous):

just in case

OpenStudy (anonymous):

you were wondering

OpenStudy (dumbcow):

ok \[x_n = x_{n-1} - \frac{f(x)}{f'(x)}\] where \[f(x) = \ln(2x) - 3^{2-x}\] the derivative \[f'(x) = \frac{1}{x} + \ln(3) 3^{2-x}\] start with an initial value .... x_0 = 1 repeat until f(x) approaches 0

OpenStudy (dumbcow):

using an excel spreadsheet will make it easier and faster to compute all the iterations

OpenStudy (anonymous):

http://www.wolframalpha.com/input/?i=ln%282x%29%3D3%5E%282-x%29 here the answer, I don't know how to solve step by step

OpenStudy (dumbcow):

thats what i was trying to do... you cant solve but you can approximate i wrote out newtons method for approximating zeroes

OpenStudy (anonymous):

thank

OpenStudy (dumbcow):

https://en.wikipedia.org/wiki/Newton%27s_method

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

thanks, i go sleep now I'm tired

OpenStudy (dumbcow):

haha ok your welcome